# Memory in the Burst Occurrence of Repeating FRBs

### v1

@@ Abstract
Understanding the nature of repeating FRBs is crucial to probe the physics of FRBs. In this work, we analyze the statistics of waiting time between bursts of three repeating FRBs from four data sets. We find a universal pronounced dependency of the waiting times on the previous time interval (denoted as $\lambda_{0}$). We observe a temporal clustering where short waiting times tend to be followed by short ones, and long by long. This memory dependency is manifested in the conditional mean waiting time as well as in the conditional mean residual time to the next burst, both of which increase in direct proportion to $\lambda_{0}$. Consequently, the likelihood of experiencing a subsequent FRB burst within a given time window after the preceding burst is significantly influenced by the burst history. We reveal that, for the first time, these memory effects are present in the scale-invariant preconditioned waiting time distribution. We show that the memory effect provides a unified description of waiting times which may account for both the repeating FRBs and the apparent non-repeating FRBs (i.e. only observed one time). These results shed new light on the mechanism of FRBs.
@@ 1 Introduction
Fast Radio Bursts (FRBs) represent one of the most powerful astronomical phenomena discovered in recent years (Cordes & Chatterjee, 2019; Bailes, 2022). Characterized by their brief yet energetic emissions, lasting merely a few milliseconds, they release energy comparable to what the sun emits over several days. The study of FRBs holds the promise of unveiling critical insights into the cosmic medium, the distribution of matter among galaxies and galaxy clusters (Petroff et al., 2019; Akahori et al., 2016). Moreover, observing the delay and frequency drift in FRBs offers valuable insights into the nature and origin of these mysterious cosmic phenomena (Hessels et al., 2019; Metzger et al., 2019), which can help constrain theoretical models and shed light on the mechanisms responsible for their generation. The research in FRBs also propels us into the realm of exploring unknown physical phenomena, including supernova explosions, neutron star mergers, black hole activities (Cordes & Chatterjee, 2019), and even exotic possibilities like axion miniclusters (Rosa & Kephart, 2018). These investigations bear significant physical implications and contribute to our understanding of fundamental astrophysical processes. One of the most significant discoveries in the study of FRBs is the identification of a repeating FRB known as FRB 121102 (Spitler et al., 2016). This repeating burst has provided valuable insights into the physical nature of FRBs and has supported models that propose an extragalactic neutron star as the origin of these bursts.
@@
The origin and propagation mechanisms surrounding FRBs, however, remain mysterious, inspiring a abundance of theories. Among the proposed models, the collision or merger of compact objects like neutron stars or black holes, and the activity of highly magnetized neutron stars, known as magnetars, have attracted substantial attention. The discovery of X-ray emissions accompanying an FRB from a magnetar in the Milky Way, for instance, hints at the plausible role of magnetars in FRB production (Bochenek et al., 2020; Zhang, 2020; Li et al., 2021a). Furthermore, the young, highly magnetized, extragalactic neutron star model, substantiated by the observation of a repeating FRB known as FRB 121102, has also found support within possible models (Spitler et al., 2016).
Despite the progress made in understanding their physical properties through the study of host galaxies, the existing models do not fully explain all aspects of FRBs (Zhang, 2020, 2023). Particularly, the phenomenon of repeating bursts - a scenario where multiple bursts emerge from the same source - poses a substantial enigma. The speculation around the nature of repeating bursts suggests a contrast; while some FRBs exhibit repeating behavior, others appear as one-off events, with no subsequent activity detected. The presence of non-repeating FRBs cannot be entirely ruled out, yet the majority of observed non-repetitive bursts may likely still be repetitive bursts (Zhang, 2023). The current unseen apparent non-repeating reasons, along with the intermittent phenomena and extended time intervals between these FRBs as exhibited in existing records, should require further investigation.
@@
Since the identification of the first FRB in 2007 (Lorimer et al., 2007), the tally of discovered FRBs has increased to over 700, with more than 60 shown repeating burst phenomena (Xu et al., 2023). The ongoing debate on the fundamental difference, if any, between the origins and physical mechanisms of repeating and non-repeating bursts continues to stimulate investigation. Recent efforts by FAST, involving continuous observations of several repeating FRBs with low threshold and high time resolution, have collected a wealth of data on repeating FRBs. This rich repository of information provides an opportunity to uncover more clues about the origin mechanisms of FRBs and the key temporal characteristics of repeating bursts through rigorous statistical analysis.
@@
The temporal analysis of repeating FRBs suggests an evolving nature of their sources over time. If the emission of FRBs followed a Poisson process, the occurrence of bursts would be random, with a consistent number within any given time frame. Nevertheless, recent research has considered the likelihood of periodicity in the timing of FRB events (The CHIME/FRB Collaboration et al., 2022; Wei et al., 2022), and evidence points to a non-Poisson distribution in some repeating FRBs (Oppermann et al., 2018; Wang et al., 2023; Du et al., 2023). A newly identified characteristic, demonstrated by the scaling properties in burst timing across diverse sources, persists despite the variable environments of these sources (Du et al., 2023). The uniformity observed in time interval distributions from different FRB sources, even when influenced by varying fluences, emphasizes the significance of inter-event time statistics in understanding the dynamics of bursts. Contrasting with the distribution of burst energy, these findings could underline the potential to uncover key dynamics of the burst mechanism through statistical examination of the event time distributions (Kumar et al., 2020).
@@
The specific distribution function, scaled by the average interval time of bursts $\bar{\lambda}$, is expressed as
@@
where $\beta$ and $\gamma$ are fitting parameters.
Based on the previously established scale-invariant waiting time distribution Eq. (1) (Du et al., 2023), this study explores further into the temporal characteristics of FRBs occurrences contained in $P$ (or $f$). Specifically, we examine the distribution of time intervals between successive bursts in repeating FRBs when the intervals between the last two bursts are $\lambda_{0}$, effectively the conditional probability distribution $P_{I}(\lambda|\lambda_{0})$. For Poisson processes, which lack memory effects, the outcome of $P_{I}(\lambda|\lambda_{0})$ is identical to $P_{I}(\lambda)$ and is independent of the value of $\lambda_{0}$. Our analysis reveals that $P_{I}(\lambda|\lambda_{0})$ is strongly dependent on the $\lambda_{0}$; smaller time intervals are more likely to be followed by smaller time intervals, and vice versa for larger time intervals. This clustering behavior in repeating FRBs is further quantified in various aspects throughout this paper, such as the average time intervals and residual time under the condition $\lambda_{0}$. Moreover, if one considers a scenario where the intervals between two consecutive bursts are $\lambda_{0}$, the results indicate that the memory effects in repeating FRBs significantly influence the probability of the next burst occurring within time $t$. This is crucial for assessing the event time of future bursts based on existing FRB observations.
2 Results
@@
For the purpose of studying the issue of temporal clustering of repeating FRBs, we search four high-quality datasets, The first dataset is FRB 121102, comprising $1652$ bursts detected over $59.5$ hours across $47$ days (Li et al., 2021b). The second dataset focuses on FRB 20201124A, featuring $1863$ bursts observed over $82$ hours spanning $54$ days (Xu et al., 2022). The third dataset is related to FRB 20220912A, which include $1076$ bursts in $8.67$ hours over approximately $55$ days(Zhang et al., 2023). The fourth dataset involves another extremely active episode of FRB 20201124A, containing $881$ bursts observed in roughly $4$ hours within a $4$-day period(Zhou et al., 2022). For the purpose of differentiation, in this paper, the two datasets related to FRB 20201124A in sequence of time will be denoted as FRB 20201124A(A) and FRB 20201124A(B) in the following sections. All datasets were acquired by FAST recently and the high resolution detection has effectively reduced the flux threshold by up to three times compared to previous observations. Recent research on FRB 121102 and FRB 20201124A has revealed an absence of periodicity or quasi-periodicity on timescales ranging from $1$ ms to $1,000$ s (Li et al., 2021b; Xu et al., 2022). In this analysis, we investigate both waiting time and its scale invariance across different fluence thresholds. We uncover novel scaling and universality phenomena previously unreported in the field.
@@
Due to the discontinuous nature of observations conducted by FAST, we redefine the waiting time as the recurrence interval between consecutive events observed within the same observational period, avoiding long gaps of observation [see Fig. 1(a)]. In our analysis, we concentrate on waiting times, denoted as $\lambda$, for bursts with fluence exceeding a specified threshold value $I$. Here, $\lambda=t_{i+1}-t_{i}$, where $t_{i}$ and $t_{i+1}$ represent the arrival times of the $i$th and $(i+1)$th burst, respectively.
@@
As a typical example, the sequence of waiting times for FRB 121102 is shown in Fig. 1(b), where the horizontal line is the median waiting time. Pronounced clusters of both short and long waiting times, relative to the median, are evident. In contrast, Fig. 1(c) displays the sequence of waiting times after shuffling the original sequence, and it notably lacks the similar clustering behavior observed in Fig. 1(b). The clusters in Fig. 1(b) qualitatively illustrate the presence of a memory effect in the waiting time sequence, where longer waiting times typically follow longer ones, and shorter times follow shorter ones.
@@
Figure 1: (a) The FAST telescope observed the sequence of FRB 121102 with a fluence greater than 0.2 Jy ms during continuous observation sessions.(b) Sequence of the waiting time above and below the median $\lambda$ = 120 s. (c) Same as (b) but for shuffled sequence of original data.
@@
To quantitatively characterize the memory effect, we examine the distribution of subsequent waiting times under different preceding conditions, i.e., different values of the previous waiting time denoted as $\lambda_{0}$, represented as $P_{I}(\lambda|\lambda_{0})$. To distinguish the magnitude of $\lambda_{0}$, the entire waiting time sequence is sorted in ascending order and divided into four subsets $q_{1},q_{2},q_{3}$ and $q_{4}$, each containing an equal number of elements. Fig. 2 provides a comparative analysis of the waiting time distribution $\bar{\lambda}P_{I}(\lambda|\lambda_{0})$ when the preceding condition $\lambda_{0}$ is placed in $q_{1}$ (right green triangle) and $q_{4}$ (left blue triangle). Additionally, the distribution of waiting times without considering the preceding condition, $\bar{\lambda}P(\lambda)$, is plotted with the complete data sequence designated as $q_{0}$ (red dots). Moreover, different fluence thresholds were selected for different FRB sources, and the actual results are independent of the magnitude of the fluence threshold. It is worth noting that a secondary peak in the waiting time, ranging from a few ms to tens of ms, was observed in all four burst source datasets, potentially due to the specific criteria used to characterize separate outbursts.(Li et al., 2021b; Jahns et al., 2023; Xu et al., 2022; Zhang et al., 2023; Zhou et al., 2022). Unlike the first peak, the location of the secondary peak remains constant as the fluences increase, disrupting the universal scaling relation (Li et al., 2021b). For the occurrence of the first peak is depend on the intensity of activity from the originating source and to ensure a uniform treatment across all distributions, we only consider the waiting times within the scaling region; that is, the minimal waiting time interval surpasses the secondary peak range, approximately $10$s for FRB 121102 and 20201124A(A), and $1$s for FRB 20220912A and 20201124A(B).
@@
Figure 2: The preconditioned distribution $P(\lambda|\lambda_{0})$ characterizes the waiting times $\lambda$ between successive FRBs that exceed specified thresholds for various sources. The preceding condition values $\lambda_{0}$ are taken from both the first (minimal) and last (maximum) quartiles, denoted as $q_{1}$ and $q_{4}$, respectively. For the sake of uniform notation, $q_{0}$ is used to represent the entire dataset, implying that $P(\lambda|\lambda_{0}\in q_{0})=P(\lambda)$. (a) $P(\lambda|\lambda_{0})$ for FRB 121102 with fluence thresholds of 0, 0.028 and 0.1 Jy ms; (b) $P(\lambda|\lambda_{0})$ for FRB 20201124A(A) with fluence thresolds of 0, 0.5, 1.0 Jy s; (c) $P(\lambda|\lambda_{0})$ for FRB 20220912A with fluence thresholds of 0, 0.1, 0.6 Jy ms; (d) $P(\lambda|\lambda_{0})$ for FRB 20201124A(B) with fluence thresholds of 0, 0.1, 0.5 Jy ms.
@@
As illustrated in Fig. 2, $P_{I}(\lambda|\lambda_{0})$ exhibits a strong dependence on the preceding condition values $\lambda_{0}$, and changes its functional form accordingly. Similar to $P_{I}(\lambda)$, $P_{I}(\lambda|\lambda_{0})$ demonstrates independence from the magnitude of fluence $I$. Hence, in subsequent analyses, the magnitude of fluence $I$ will be omitted unless explicitly needed. As observed from Fig. 2, when $\lambda_{0}$ is taken from $q_{1}$ and $q_{4}$ respectively, shorter waiting times are more probable ($P(\lambda|\lambda_{0})$) in the interval $\lambda/\bar{\lambda}<1$ compared to longer waiting times. Conversely, for $\lambda/\bar{\lambda}>1$, the situation is reversed. Across the entire range, the $P(\lambda|\lambda_{0})$ without any preceding condition ($\lambda_{0}\in q_{0}$) falls between theses two cases. Further analysis reveals that $P(\lambda|\lambda_{0})$ can also be well approximated by $P(\lambda|\lambda_{0})\sim f(\lambda/\bar{\lambda},\lambda_{0}/\bar{\lambda})/\bar{\lambda}$. Deviations from an exponential function further demonstrates the strong memory effects that the preceding condition values $\lambda_{0}$ has on subsequent FRB bursts and their degree of clustering. The parameter $\gamma$ in Eq. (2) describes these effects, with more details provided in Table 1.

### v2

@@ Abstract
Understanding the nature of repeating fast radio bursts (FRBs) is crucial to probe their underlying physics. In this work, we analyze the waiting time statistics between bursts of three repeating FRBs from four datasets. We find a universally pronounced dependency of the waiting times on the previous time interval (denoted as $\lambda_{0}$). We observe a temporal clustering where short waiting times tend to be followed by short ones, and long by long comparative to their mean value. This memory dependency is manifested in the conditional mean waiting time as well as in the conditional mean residual time to the next burst, both of which increase in direct proportion to $\lambda_{0}$. Consequently, the likelihood of experiencing a subsequent FRB burst within a given time window after the preceding burst is generally influenced by the burst history. We reveal that for the first time, these memory effects are present in the scale-invariant preconditioned waiting time distribution. We show that the memory effect provides a unified description of waiting times which may account for both the repeating FRBs and the apparent non-repeating FRBs (i.e. only observed one time). These results shed new light on the mechanism of FRBs.
@@ 1 Introduction
Fast Radio Bursts (FRBs) represent one of the most powerful astronomical phenomena discovered in recent years (Cordes & Chatterjee, 2019; Bailes, 2022). Characterized by their brief yet energetic emissions, lasting merely a few milliseconds, they release energy comparable to what the sun emits over several days. The study of FRBs holds the promise of unveiling critical insights into the cosmic medium, the distribution of matter among galaxies and galaxy clusters (Akahori et al., 2016; Petroff et al., 2019). Moreover, observing the delay and frequency drift in FRBs offers valuable insights into the nature and origin of these mysterious cosmic phenomena (Hessels et al., 2019; Metzger et al., 2019), which can help constrain theoretical models and shed light on the mechanisms responsible for their generation. These investigations bear significant physical implications and contribute to our understanding of fundamental astrophysical processes. One of the most significant discoveries in the study of FRBs is the identification of a repeating FRB known as FRB 20121102A (Spitler et al., 2016). This repeating burst has provided valuable insights into the physical nature of FRBs and has supported models that propose an extragalactic neutron star as the origin of these bursts (Hessels, 2017; Josephy et al., 2019; Lyu et al., 2021; Li et al., 2022).
@@
However, the origin and propagation mechanisms surrounding FRBs remain mysterious, inspiring an abundance of theories. Among the proposed models, the collision or merger of compact objects like neutron stars or black holes, and the activity of highly magnetized neutron stars, known as magnetars, have attracted substantial attention. The discovery of X-ray emissions accompanying an FRB from a magnetar in the Milky Way, for instance, hints at the plausible role of magnetars in FRB production (Bochenek et al., 2020; Andersen et al., 2020; Li et al., 2021a). Furthermore, the young, highly magnetized, extragalactic neutron star model, supported by the observation of FRB 20121102A, has also found support within possible models (Spitler et al., 2016).
Despite the progress made in understanding their physical properties through the study of host galaxies, current models do not fully explain all aspects of FRBs (Zhang, 2020, 2023; Kirsten et al., 2024). In particular, the phenomenon of repeating bursts, in which multiple bursts emerge from the same source, remains a substantial enigma. Although some FRBs are observed to repeat, others appear as one-off events without subsequent detections. This contrast raises intriguing questions about the nature of these bursts. The presence of non-repeating FRBs is clear; however, it remains an open question whether these observed non-repetitive bursts might still exhibit repetitive behavior under extended observation periods or with improved detection technologies. As noted in Zhang (2023), the hypothesis that all FRBs could potentially repeat has not been definitively ruled out and requires further investigation. Supporting this notion, Kirsten et al. (2024) provide empirical evidence suggesting that high-energy burst distributions of certain hyperactive repeating sources resemble those of the non-repeating FRB population. This observation indicates that apparently non-repeating FRB sources may simply be the rarest bursts from repeating sources, which could be observable across different energy scales. Continued research into the intermittent phenomena and the extended time intervals between these FRBs, as exhibited in existing records, is essential to develop a more complete understanding of FRB activity. The findings of Kirsten et al. (2024) underscore the complexity of categorizing FRBs definitively as repeaters or non-repeaters based on current observational capabilities and highlight the need for sustained observational efforts across multiple energy thresholds.
@@
Since the identification of the first FRB in 2007 (Lorimer et al., 2007), the tally of discovered FRBs has increased to more than 700, showing over 60 instances of repeated burst phenomena (Xu et al., 2023). The ongoing debate on the fundamental difference, if any, between the origins and physical mechanisms of repeating and non-repeating bursts continues to stimulate investigation. Recent efforts by FAST, involving continuous observations of several repeating FRBs with low threshold and high time resolution, have collected a wealth of data on repeating FRBs. This rich repository of information provides an opportunity to uncover more clues about the origin mechanisms of FRBs and the key temporal characteristics of repeating bursts through rigorous statistical analysis.
@@
The temporal analysis of repeating FRBs suggests an evolving nature of their sources over time. If the emission of FRBs followed a Poisson process, the occurrence of bursts would be random, with a consistent number within any given time frame. However, recent research has considered the likelihood of periodicity in the timing of FRB events (The CHIME/FRB Collaboration et al., 2022; Wei et al., 2022), and evidence points to a non-Poisson distribution in some repeating FRBs (Wang & Yu, 2017; Oppermann et al., 2018; Wang et al., 2023; Du et al., 2024). A newly identified characteristic, demonstrated by the scaling properties in burst timing across diverse sources, persists despite the variable environments of these sources (Du et al., 2024). The uniformity observed in time interval distributions from different FRB sources, even when influenced by varying fluences, emphasizes the significance of interevent time statistics in understanding the dynamics of bursts. Contrasting with the distribution of burst energy, these findings could underline the potential to uncover key dynamics of the burst mechanism through statistical examination of the event time distributions (Kumar et al., 2020).
@@
The specific distribution function, scaled by the average interval time of the bursts $\bar{\lambda}$, is expressed as
@@
where $\beta$ and $\gamma$ are fitting parameters. Specifically, for $\gamma$ as the shape parameter, which determines how the tail of the distribution decays. The Pareto distribution, known for its heavy-tailed nature, is well-suited for modeling phenomena where rare but significant events have a large impact, such as the intense intermitent bursts observed in the waiting time distribution of repeating FRBs. Unlike the Weibull distribution, which is customized to model both clustered and spread-out events through its scale and shape parameters, the Pareto distribution effectively captures the sporadic nature of FRBs, where long periods of inactivity are interrupted by sudden, extreme bursts. By focusing on the occurrence of these irregular events, the Pareto distribution may provide a more fitting framework than non-Poisson distributions like the Weibull, particularly in accurately representing the extreme values denoted by heavy-tail is crucial to understanding FRBs dynamics.
Based on the previously established scale-invariant waiting time distribution Eq. 1, $P$ (or $f$) (Du et al., 2024), this study investigates the temporal characteristics of the occurrences of FRB, specifically focusing on the distribution of time intervals between successive bursts in repeating FRBs. We analyze the conditional probability distribution $P_{I}(\lambda|\lambda_{0})$, where $\lambda_{0}$ represents the interval between the last two bursts. For Poisson processes, which lack memory effects, $P_{I}(\lambda|\lambda_{0})$ is expected to be identical to $P_{I}(\lambda)$ and independent of $\lambda_{0}$. However, our study aims to explore potential deviations from this behavior in repeating FRBs, which may indicate the presence of memory effects. We further quantify these temporal characteristics through various statistical measures, such as average time intervals and residual times under different conditions. These findings could provide insights into the probability of future burst occurrences based on existing observations of FRBs, contributing to a deeper understanding of the mechanisms governing FRB activity. Additionally, we propose a unified framework that might explain the waiting times of both repeating and apparently non-repeating FRBs.
2 Method and Results
@@
For the purpose of studying the issue of temporal clustering of repeating FRBs, we search four high-quality datasets, The first dataset is FRB 20121102A, comprising $1652$ bursts detected over $59.5$ hours across $47$ days (Li et al., 2021b). The second dataset focuses on FRB 20201124A, featuring $1863$ bursts observed over $82$ hours spanning $54$ days (Xu et al., 2022). The third dataset is related to FRB 20220912A, which includes $1076$ bursts in $8.67$ hours over approximately $55$ days (Zhang et al., 2023). The fourth dataset involves another extremely active episode of FRB 20201124A, containing $881$ bursts observed in approximately $4$ hours within a $4$-day period (Zhou et al., 2022). For the purpose of differentiation, in this paper, the two datasets related to FRB 20201124A in sequence of time will be denoted as FRB 20201124A(A) and FRB 20201124A(B) in the following sections. All datasets were recently acquired by FAST, which provided high raw sensitivity and temporal resolution detection (Qian et al., 2020), effectively reducing the flux threshold by up to three times compared to previous observations for FRB 20121102A (refer to Li et al. (2021b) and the references cited therein). Recent research on FRB 20121102A and FRB 20201124A has revealed an absence of periodicity or quasiperiodicity on timescales ranging from $1$ ms to $1,000$ s (Li et al., 2021b; Xu et al., 2022).
@@
Due to the discontinuous nature of the observations conducted by FAST, we redefine the waiting time as the recurrence interval between consecutive events observed within the same observational period [see Fig. 1(a)]. In our analysis, we focus on waiting times, denoted $\lambda$, for bursts with fluence exceeding a specified threshold value $I$. Here, $\lambda=t_{i+1}-t_{i}$, where $t_{i}$ and $t_{i+1}$ represent the arrival times of the $i$ th and $(i+1)$ th bursts, respectively. Throughout our statistical calculations, we use logarithmic binning to ensure that each timescale is appropriately represented. Despite this, the calculation of the probability density function (PDF) makes the exact number and width of the bins less significant. In this context, the uncertainty associated with the counts of waiting times is estimated by $\sqrt{N_{j}}$, where $N_{j}$ represents the number of waiting times counted in the jth bin (Lin & Sang, 2019; Wei et al., 2021).
@@
As a typical example, the sequence of waiting times for FRB 20121102A is shown in Fig. 1(b), where the horizontal line is the median waiting time. The pattern of clusters of short and long waiting times, relative to the median, is apparent. In contrast, Fig. 1(c) displays the sequence of waiting times after shuffling the original sequence, and it notably lacks the similar clustering behavior observed in Fig. 1(b). The clusters in Fig. 1(b) qualitatively illustrate the presence of a memory effect in the waiting time sequence, where longer waiting times typically follow longer ones and shorter times follow shorter ones.
@@
Figure 1: (a) The FAST telescope observed the sequence of FRB 121102 with a fluence greater than 0.2 Jy ms during continuous observation sessions.(b) Sequence of the waiting time above and below the median $\lambda$ = 120 s (red line). (c) Same as (b) but for shuffled sequence of original data.
To quantitatively characterize the memory effect, we examine the distribution of subsequent waiting times under different preceding conditions, that is, different values of the previous waiting time denoted as $\lambda_{0}$, represented as $P_{I}(\lambda|\lambda_{0})$. To distinguish the magnitude of $\lambda_{0}$, the entire waiting time sequence is sorted in ascending order in time and divided into four subsets $q_{1},q_{2},q_{3}$ and $q_{4}$, each containing an equal number of elements. Fig. 2 provides a comparative analysis of the waiting time distribution $\bar{\lambda}P_{I}(\lambda|\lambda_{0})$ when the preceding condition $\lambda_{0}$ is placed in $q_{1}$ (green symbols) and $q_{4}$ (blue symbols). Furthermore, the distribution of waiting times without considering the preceding condition, $\bar{\lambda}P(\lambda)$, is plotted with the complete data sequence designated as $q_{0}$ (red symbols). In addition, different fluence thresholds are shown for different sources of FRB with various symbols, and the actual results are statistically independent of the magnitude of the fluence threshold. For the same FRB source and under the same preconditions, it can be observed that the waiting times have nearly the same distribution form under different fluence thresholds. The best-fit lines (the dashed lines) are obtained using the Markov Chain Monte Carlo (MCMC) method, and more detailed fitting information is listed in Table 1. From the table, we observe differences in the $\gamma$ values between certain datasets, especially for datasets with preconditions from $q_{1}$ and $q_{4}$, which exhibit a statistical significance approximately from $1.0$ to $3.5\sigma$ (3.5, 0.94, 2.0, 3.2 for FRB 20121102A, FRB 20220912A, FRB 20201124A(A) and FRB 20201124A(B), respectively). This indicates a moderate to strong likelihood that the observed differences are not due to random chance, suggesting potential underlying variations between these comparisons of $q_{1}$ and $q_{4}$. However, the results of $0.94\sigma$ and $2.0\sigma$ are less significant and indicate that, for these particular comparisons, the evidence for the differences is weaker and could more likely be due to statistical fluctuations for limited samples. These mixed significance levels highlight the complexity of interpreting results with the current dataset size and the variability inherent in FRBs. Although some findings point to meaningful differences, they fall short of the more rigorous thresholds typically required for definitive conclusions. This underscores the need for caution when interpreting the results and highlights the importance of acquiring more extensive data. Furthermore, we compared the significance of the parameter $\gamma$ under different preconditions $(q_{0},q_{1},q_{2},q_{3},q_{4})$ for FRB 20201124A(A) and FRB 20201124A(B). The significance levels were found to be $(0.59,0.39,0.19,0.41,0.80)$, indicating similarity within $1\sigma$. Given that FRB 20201124A(B) occurred during a highly active period spanning 4 days, this similarity in $\gamma$ suggests an intrinsic connection in the burst behavior of the repeating FRB source during different active periods. This finding further supports the idea that the parameter $\gamma$ may capture underlying mechanisms that govern FRB activity, even across varying levels of source activity.
@@
It is worth noting that a secondary peak in waiting time, ranging from a few to tens of milliseconds, was observed in the four burst source datasets, possibly due to the specific criteria used to characterize separate outbursts (Li et al., 2021b; Jahns et al., 2023; Xu et al., 2022; Zhang et al., 2023; Zhou et al., 2022). Unlike the first peak, the location of the secondary peak remains constant as the fluence increases (Li et al., 2021b), disrupting the universal scaling relation (Du et al., 2024). As the occurrence of the first peak depends on the intensity of activity from the originating source and to ensure a uniform treatment across all distributions, we only consider the waiting times within the scaling region; that is, the minimal waiting time interval surpasses the secondary peak range, approximately $10$ s for FRB 20121102A and FRB 20201124A(A), and $1$ s for FRB 20220912A and FRB 20201124A(B).
@@
Figure 2: The preconditioned distribution $P(\lambda|\lambda_{0})$ characterizes the waiting times $\lambda$ between successive FRBs that exceed specified thresholds for various sources. The preceding condition values $\lambda_{0}$ are taken from both the first (minimal) and last (maximum) quartiles, denoted as $q_{1}$ and $q_{4}$, respectively. For the sake of uniform notation, $q_{0}$ is used to represent the entire dataset, implying that $P(\lambda|\lambda_{0}\in q_{0})=P(\lambda)$. (a) $P(\lambda|\lambda_{0})$ for FRB 20121102A with fluence thresholds of 0.0 (square), 0.028 (circle) and 0.1 (right triangle) Jy ms; (b) $P(\lambda|\lambda_{0})$ for FRB 20201124A(A) with fluence thresholds of 0.0 (square), 0.5 (circle), 1.0 (right triangle) Jy s; (c) $P(\lambda|\lambda_{0})$ for FRB 20220912A with fluence thresholds of 0.0 (square), 0.1 (circle), 0.6 (right triangle) Jy ms; (d) $P(\lambda|\lambda_{0})$ for FRB 20201124A(B) with fluence thresholds of 0.0 (square), 0.1 (circle), 0.5 (right triangle) Jy ms. For all subplots, the colors red, green, and blue correspond to $\lambda_{0}$ belonging to $q_{0}$, $q_{1}$, and $q_{4}$, respectively. The insets illustrate the behavior of the fitting curves near the point where $\lambda/\bar{\lambda}=1$. The residuals between the data points and the best-fit values relative to the errors for all data are also displayed, with the color scheme corresponding to the different preconditions consistent with those in the preconditioned distribution $P(\lambda|\lambda_{0})$.
@@
As illustrated in Fig. 2, $P_{I}(\lambda|\lambda_{0})$ generally exhibits an evident dependence on the preceding condition values $\lambda_{0}$, and changes its functional form accordingly. Later, when performing the specific calculations for a two-sample Kolmogorov-Smirnov (K-S) test, we will quantitatively discuss the difference of the distributions $P_{I}(\lambda|\lambda_{0})$ again. Similarly to $P_{I}(\lambda)$, $P_{I}(\lambda|\lambda_{0})$ demonstrates independence from the magnitude of fluence $I$. Hence, in subsequent analyzes, the magnitude of fluence $I$ will be omitted unless explicitly needed. As observed in Fig. 2, when $\lambda_{0}$ is taken from $q_{1}$ and $q_{4}$, respectively, shorter waiting times are more probable ($P(\lambda|\lambda_{0})$) in the interval $\lambda/\bar{\lambda}<1$ compared to longer waiting times. In contrast, for $\lambda/\bar{\lambda}>1$, the situation is reversed. Across the entire range, the $P(\lambda|\lambda_{0})$ without any preceding condition ($\lambda_{0}\in q_{0}$) falls between these two cases.
Further fitting analysis reveals that $P(\lambda|\lambda_{0})$ can also be well approximated by $P(\lambda|\lambda_{0})\sim f(\lambda/\bar{\lambda},\lambda_{0}/\bar{\lambda})/\bar{\lambda}$. Deviations from an exponential function can be well described by the Tsallis q-exponential function (Du et al., 2024). The parameter $\gamma$ in Eq. (2) plays a key role in demonstrating the memory effects that the preceding condition values $\lambda_{0}$ have on subsequent FRB bursts and their degree of clustering. More details about $\gamma$ that are affected by different preceding conditions are provided in Table 1. For preconditioned intervals from $q_{1}$, $q_{2}$, $q_{3}$ to $q_{4}$, the parameter $\gamma$ generally decreases monotonically, a pattern that holds across the four datasets. However, considering the presence of fitting errors, there is a possible likelihood that overlapping intervals for $\gamma$ could occur between adjacent preconditioned intervals, such as between $[q_{1},q_{2}]$, $[q_{2},q_{3}]$, and $[q_{3},q_{4}]$, and we also cannot completely rule out the possibility of overlapping intervals between $[q_{1},q_{4}]$. After all, current statistical data samples are limited, and grouping further reduces the number of samples, which may lead to fluctuations in the statistical results. Furthermore, if the uncertainty of waiting times counts in the time intervals is considered, the parameter $\gamma$ may differ compared to scenarios where this uncertainty is not taken into account. However, a precise determination of the values of the model parameter is not the main objective of this study. Instead, we primarily explore how different preconditioning, such as $q_{1}$, $q_{2}$, $q_{3}$, or $q_{4}$, influences the time intervals between subsequent FRB bursts, essentially studying the memory effect.